# From RNNs to Foundation Models: An Empirical Study on Commercial Building Energy Consumption

### v1

@@ Appendix D Hyperparameter Tuning
Given these models span a wide range of architectures, using the same value for the learning rate for all of them would lead to poor performance. To that end, we tuned the learning rate over the validation set of the problem $(T=96,L=48)$ with choices $\texttt{lr}=\{10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5},10^{-5},5\times 10^{-6}\}$. The results are reported in Table 8.
Table 8: Learning rate tuning over the validation set of the problem $(L=96,T=48)$. The metric is MSE, and the best learning rates are highlighted.
Model $10^{-3}$ $5\times 10^{-4}$ $10^{-4}$ $5\times 10^{-5}$ $10^{-5}$ $5\times 10^{-6}$
LSTM 2.3091 5.3035 0.4369 0.3611 0.3807 0.6348
LSTNet 0.9551 0.5223 0.1652 0.1246 0.1573 0.2479
Transformer 9.3107 8.7265 1.5617 0.8678 2.3310 1.2849
Autoformer 0.4911 0.3675 0.2472 0.1634 0.2441 0.3875
Informer 8.7894 10.1551 9.1948 15.2048 2.9153 0.6201
TimesNet 0.0781 0.0845 0.0807 0.0796 0.0773 0.0814
Since Informer showed the best performance at the end of the learning rate choices, we ran a further experiment to test whether even lower learning rates are better for it, whose conclusion was negative (see Table 9).
Table 9: Extended learning rate tuning for Informer.
Model $10^{-6}$ $5\times 10^{-7}$ $10^{-7}$
Informer 2.5689 4.1384 1.2238
@@
LSTM $20^{*}$ 1 $20^{*}$ $20^{*}$ $20^{*}$ $20^{*}$
LSTNet 7 15 5 19 $20^{*}$ $20^{*}$
Transformer 7 2 8 19 $20^{*}$ $20^{*}$
Autoformer 5 1 1 9 15 2
Informer 10 2 2 7 17 15
TimesNet 2 1 19 19 $20^{*}$ $20^{*}$
The observations of Table 10 can be explained by the fact that training of IL-HET is more unstable than IL-HOM, which leads to early stopping being triggered very early into the training run. This is further verified by letting Transformer for $L=96,T=96$ train to completion without early stopping on IL-HET, which reveals instability in training around epoch 8 (see Figure 4).
Figure 4: Training instability in Transformer training on IL-HET.
Similar to the base models, the learning rate for TimeFM was tuned over the validation set of the problem $(T=96,L=96)$ with choices $\texttt{lr}=\{10^{-2},10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5},10^{-5},5\times 10^{-6}\}$ over 1 epoch. The results are reported in Table 11. Note that $10^{-2}$ is the default learning rate in the pre-trained TimesFM model.
Table 11: Learning rate tuning over the validation set of the problem $(L=96,T=96)$. The metric is MSE, and the best learning rates are highlighted.
Model $10^{-2}$ $10^{-3}$ $5\times 10^{-4}$ $10^{-4}$ $5\times 10^{-5}$ $10^{-5}$ $5\times 10^{-6}$
TimesFM 0.2034 0.4440 0.5235 0.6985 0.7347 0.7566 0.7582
Patience is also implemented in fine-tuning the TimesFM model where validation is performed to compute the progress of the fine-tuning in terms of the MSE loss every 1,000 steps in each epoch. Patience is increased by one if the MSE loss does not decrease and the so-called early termination of the current epoch is executed if patience reaches five. In our fine-tuning experience, the early termination condition was never triggered.
@@ Appendix E Predictions
Figure 5: Predictions of the test set loads for two selected clients from IL-HOM and IL-HET datasets, respectively, for the best models reported in Table 6.

### v2

@@ Appendix D Hyperparameter Tuning
Given these models span a wide range of architectures, using the same value for the learning rate for all of them would lead to poor performance. To that end, we tuned the learning rate over the validation set for all lookaheads ($T=4,48,96$) and all datasets (IL-HET and IL-HOM) with choices $\texttt{lr}=\{10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5},10^{-5},5\times 10^{-6},10^{-6},5\times 10^{-7}\}$. The results are reported in Tables 8 and 9.
Table 8: Learning rate tuning over the validation set on the dataset IL-HET. The metric is NMSE, and the best learning rates are highlighted.
Model $10^{-3}$ $5\times 10^{-4}$ $10^{-4}$ $5\times 10^{-5}$ $10^{-5}$ $5\times 10^{-6}$ $10^{-6}$ $5\times 10^{-7}$
$L=512,T=4$
LSTM 0.8258 0.2535 0.2020 0.1853 0.2804 0.2823 0.3878 0.4077
LSTNet 0.4564 0.2236 0.1417 0.1709 0.2202 0.3225 0.3207 0.3397
Transformer 3.0464 3.0721 0.8025 1.2834 1.5921 1.0574 0.4246 0.6596
Autoformer 0.6904 0.5245 0.5835 0.4326 0.3083 0.3834 0.2688 0.2649
Informer 3.9926 3.7363 3.0798 3.1364 1.5351 0.7345 1.0424 1.4432
TimesNet 0.0699 0.0667 0.0728 0.0627 0.0603 0.0703 0.1000 0.1088
PatchTST 0.0678 0.0622 0.0628 0.0597 0.0619 0.0666 0.0829 0.0924
$L=512,T=48$
LSTM 3.8491 2.1324 0.3300 0.2981 0.3283 0.3052 0.3924 0.4104
LSTNet 0.2866 0.3833 0.2166 0.2125 0.3213 0.3544 0.3315 0.3432
Transformer 3.3272 3.3638 0.7665 0.8136 1.6502 1.0295 0.4785 0.7056
Autoformer 0.6104 0.5294 0.5150 0.4226 0.4226 0.3457 0.2744 0.2761
Informer 2.8727 3.0340 3.3786 3.6050 1.2244 0.3451 1.4153 1.9887
TimesNet 0.1043 0.1096 0.1117 0.1032 0.0976 0.0985 0.1121 0.1175
PatchTST 0.0967 0.0953 0.1009 0.0955 0.0984 0.1034 0.1082 0.1122
$L=512,T=96$
LSTM 0.7134 0.3811 0.2001 0.1843 0.2719 0.2809 0.3868 0.4069
LSTNet 0.8501 0.2931 0.1617 0.1595 0.2360 0.3235 0.3202 0.3397
Transformer 2.7635 3.0211 0.5371 1.3160 1.7465 1.1508 0.4568 0.6904
Autoformer 0.7147 0.4788 0.5125 0.4777 0.3339 0.3632 0.2721 0.2704
Informer 3.2822 3.1977 2.9025 3.6369 1.2768 1.4782 1.1671 2.0538
TimesNet 0.0966 0.0931 0.0973 0.0914 0.0891 0.0985 0.1129 0.1203
PatchTST 0.0874 0.0874 0.0897 0.0876 0.0908 0.0951 0.1014 0.1053
Table 9: Learning rate tuning over the validation set on the dataset IL-HOM. The metric is NMSE, and the best learning rates are highlighted.
Model $10^{-3}$ $5\times 10^{-4}$ $10^{-4}$ $5\times 10^{-5}$ $10^{-5}$ $5\times 10^{-6}$ $10^{-6}$ $5\times 10^{-7}$
$L=512,T=4$
LSTM 0.2058 0.1832 0.1554 0.1618 0.2546 0.3488 0.5317 0.5421
LSTNet 0.2145 0.1604 0.1490 0.1819 0.3139 0.3417 0.4325 0.4871
Transformer 0.6288 0.5218 0.2916 0.2189 0.2717 0.2736 0.3208 0.3363
Autoformer 0.3000 0.2691 0.2050 0.2067 0.2350 0.2452 0.3032 0.3404
Informer 0.6004 0.5635 0.3656 0.2733 0.2078 0.2143 0.3040 0.3499
TimesNet 0.0835 0.0804 0.0933 0.0885 0.0976 0.1105 0.1863 0.2157
PatchTST 0.1049 0.0908 0.0987 0.1007 0.1099 0.1196 0.1466 0.1679
$L=512,T=48$
LSTM 0.3416 0.3173 0.3278 0.3371 0.3558 0.3897 0.5408 0.5500
LSTNet 0.2479 0.2456 0.2631 0.2912 0.4025 0.4197 0.4526 0.4988
Transformer 0.6229 0.3294 0.2991 0.3052 0.2859 0.2834 0.3273 0.3404
Autoformer 0.2992 0.2806 0.2806 0.2762 0.3030 0.3182 0.3428 0.3643
Informer 0.6450 0.5477 0.4009 0.3403 0.2319 0.2385 0.3393 0.3826
TimesNet 0.1439 0.1327 0.1258 0.1234 0.1322 0.1432 0.2065 0.2307
PatchTST 0.1973 0.1889 0.1994 0.1927 0.1986 0.2017 0.2176 0.2283
$L=512,T=96$
LSTM 0.2997 0.2389 0.2321 0.2402 0.3000 0.3673 0.5341 0.5444
LSTNet 0.2422 0.2350 0.2257 0.2566 0.3508 0.3741 0.4422 0.4911
Transformer 0.6371 0.3380 0.2984 0.3042 0.2915 0.2897 0.3288 0.3407
Autoformer 0.2990 0.2743 0.2401 0.2448 0.2766 0.2868 0.3327 0.3564
Informer 0.6068 0.5414 0.4018 0.3598 0.2571 0.2774 0.3366 0.3705
TimesNet 0.1503 0.1443 0.1325 0.1299 0.1523 0.1745 0.2247 0.2515
PatchTST 0.1855 0.1830 0.1856 0.1809 0.1883 0.1911 0.2037 0.2114
@@
LSTM 18 $20^{*}$ $20^{*}$ 4 4 17
LSTNet 2 12 6 10 6 8
Transformer 4 4 4 $20^{*}$ $20^{*}$ $20^{*}$
Autoformer $20^{*}$ $20^{*}$ $20^{*}$ $20^{*}$ 1 5
Informer 2 1 11 6 3 11
TimesNet 3 3 1 $20^{*}$ $7$ 15
PatchTST 1 5 3 8 11 1
The observations of Table 10 can be explained by the fact that training of IL-HET is more unstable than IL-HOM, which leads to early stopping being triggered very early into the training run. This is further verified by letting Transformer for $L=512,T=96$ train to completion without early stopping on IL-HET, which reveals instability in training after epoch 4 (see Figure 4).
Figure 4: Training instability in Transformer training on IL-HET compared to IL-HOM.
Similar to the base models, the learning rate for TimeFM was tuned over the validation set of all lookaheads $(T=4,48,96)$ with choices $\texttt{lr}=\{10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5},10^{-5},5\times 10^{-6},10^{-6},5\times 10^{-7}\}$ over 1 epoch. The results are reported in Tables 11 and 12. With $T=(48,96)$, we perform additional experiments with $\texttt{lr}=\{10^{-7},5\times 10^{-8}\}$.
Table 11: Learning rate tuning over the validation set of the problem for TimesFM. The metric is NMSE, and the best learning rates are highlighted.
Model $10^{-3}$ $5\times 10^{-4}$ $10^{-4}$ $5\times 10^{-5}$ $10^{-5}$ $5\times 10^{-6}$ $10^{-6}$ $5\times 10^{-7}$
$(L,T)$ IL-HOM
$(512,4)$ 639.06 6.7198 0.0177 0.0166 0.0074 0.0076 0.0077 0.0082
$(512,48)$ 656.45 7.7152 0.0875 0.0512 0.0384 0.0365 0.0383 0.0404
$(512,96)$ 169.77 0.9800 0.1185 0.0483 0.0403 0.0401 0.0413 0.0432
$(L,T)$ IL-HET
$(512,4)$ 2285.12 89.396 0.0331 0.0252 0.0144 0.0121 0.0118 0.0124
$(512,48)$ 247.92 10.034 0.0814 0.0724 0.0486 0.0473 0.0398 0.0369
$(512,96)$ 753.21 1.9964 0.0539 0.0568 0.0465 0.0462 0.0383 0.0369
Table 12: Additional learning rate tuning over the validation set of the problem for TimesFM for IL-HET with $(512,48)$ and $(512,96)$. The metric is NMSE, and the best learning rates are highlighted.
Model $10^{-7}$ $5\times 10^{-8}$
$(L,T)$ IL-HET
$(512,48)$ 0.04121 0.04711
$(512,96)$ 0.04027 0.04585
Early-stopping based on patience is also implemented in finetuning the TimesFM model where validation is performed to compute the progress of the finetuning in terms of the NMSE loss every 1,000 steps in each epoch. Patience is increased by one if the NMSE loss does not decrease and the so-called early termination of the current epoch is executed if patience reaches five. In our finetuning experience, the early termination condition was never triggered.
@@ Appendix E Predictions
Figure 5: Predictions of the test set loads for two selected buildings from IL-HOM and IL-HET datasets for $T=512,L=48$, respectively, for the best models reported in Table 6.